# GraphPipe: Improving Performance and Scalability of DNN Training with Graph Pipeline Parallelism

### v1

@@ 1. Introduction
First, opportunities to exploit the inherent parallel structures of a DNN are left unseized. DNN applications such as healthcare (krumholz2016data, ; tan2020multimodal, ; suresh2020deep, ), chatbot (openai2023gpt4, ), and recommendation (naumov2019deep, ) jointly process heterogeneous data types (e.g., text, images, and tabular data). DNNs employed therein are designed to feature multiple branches, which are computationally independent and thus can be executed concurrently. But existing DNN systems with SPP first linearize the computation graph of a DNN to construct the stages of a sequential pipeline and process these stages sequentially, falling short in harnessing the opportunity to blend such branch-level parallelism with pipeline parallelism.
@@ 5. Pipeline Stage Partitioner
22:function DP($\mathcal{G},c_{f},c_{b},d,t_{max}$)
23:    if this DP state has been visited then
24:         return corresponding $\mathcal{G}_{S}^{best}$ to this DP state
25:    // Consider a given DP state as a SINGLE stage
26:    $\mathcal{G}_{S}^{best}=\varnothing$
27:    if EstimateTPS($\mathcal{G},c_{f},c_{b},d$)$\leq t_{max}$ then
28:         // Optimize schedule via Algorithm 2
29:         $\Pi_{opt}$ = ScheduleStage($\mathcal{G},c_{f},c_{b},d$)
30:         $\mathcal{G}_{S}^{best}=$ StageGraph($\mathcal{G},\Pi_{opt},d$)
31:    // Decompose a given DP state into two stages
32:    if $\mathcal{G}$ can be decomposed in series then
33:         for $(\mathcal{G}_{1},\mathcal{G}_{2})\in$ SeriesDecompose($\mathcal{G}$do
34:             for $d_{2}\leftarrow 1$ to $d-1$ do
35:                 $d_{1}=d-d_{2}$
36:                 for $c_{m}\in C$ do
37:                     $\mathcal{G}_{S_{2}}^{new}=$ DP($\mathcal{G}_{2},c_{m},c_{b},d_{2},t_{max}$)
38:                     Update $i_{m}$ based on $\mathcal{G}_{S_{2}}^{new}$
39:                     $\mathcal{G}_{S_{1}}^{new}=$DP($\mathcal{G}_{1},c_{f},c_{m},d_{1},t_{max}$)
40:    else if $\mathcal{G}$ can be decomposed in parallel then
41:         for $(\mathcal{G}_{1},\mathcal{G}_{2})\in$ ParallelDecompose($\mathcal{G}$do
42:             for $d_{1}\leftarrow 1$ to $d-1$ do
43:                 $d_{2}=d-d_{1}$
44:                 $\mathcal{G}_{S_{1}}^{new}=$ DP($\mathcal{G}_{1},c_{f},c_{b},d_{1},t_{max}$)
45:                 $\mathcal{G}_{S_{2}}^{new}=$ DP($\mathcal{G}_{2},c_{f},c_{b},d_{2},t_{max}$)
46:    $\mathcal{G}_{S}^{best}=$ PickBetter($\mathcal{G}_{S}^{best},\mathcal{G}_{S_{1}}^{new}\cup\mathcal{G}_{S_{2}}^{new}$)
47:    return $\mathcal{G}_{S}^{best}$
@@
Dynamic programming (DP). GraphPipe adopts a dynamic programming algorithm where the value of each DP state indicates the existence of a strategy achieving a throughput within a target range (Lines 1320 of Algorithm 1). At each DP level, GraphPipe applies series-parallel decompositions to split an input graph (say $\mathcal{G}$) into two new subgraphs (say $\mathcal{G}_{1},\mathcal{G}_{2}$), each of which serves as the input computation graph of a new DP subproblem at one DP level below. GraphPipe recursively solves the DP subproblems to construct a solution of the original problem where the input computation graph is $\mathcal{G}_{C}$ (Lines 2247 of Algorithm 1).
@@
  • Base case: We consider the entire subgraph $\mathcal{G}$ as a single stage and apply data parallelism with a data-parallel degree of $d$ (Line 27 in Algorithm 1). We check if the target TPS range is achievable with the memory constraint, and compute the number of in-flight micro-batches according to Algorithm 2 (see §6).
  • Series decomposition: We perform a series decomposition to create two subgraphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, where the sink node of $\mathcal{G}_{1}$ coincides with the source node of $\mathcal{G}_{2}$ (Line 32 in Algorithm 1). We first solve the subproblem associated with $\mathcal{G}_{2}$. To do so, we enumerate all feasible schedules for the source node of $\mathcal{G}_{2}$. We then solve the subproblem associated with $\mathcal{G}_{1}$.
  • Parallel decomposition: We perform a parallel decomposition to create $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, where $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ share the same source and sink nodes (Line 40 in Algorithm 1). As there is no data dependency between these subgraphs, the pipelines can be executed in parallel. The subproblems associated with $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ may produce different optimal numbers of in-flight micro-batches for the shared source node. To ensure continuous pipelining, we take the larger number of in-flight micro-batches as the solution.
Overall process. Figure 4 visualizes the overall process. At the top, a DP subproblem is provided with its initial conditions: computation graph $\mathcal{G}$, the number of available devices $d$, and the target TPS range $[0,t_{max}]$. Suppose the number of in-flight micro-batches for the sink node is $i_{b}$, the micro-batch sizes for the source and sink nodes are $b_{f}$ and $b_{b}$, the stage containing the source node (i.e., source stage) uses the $k_{f}$$k_{f}$ schedule, and the stage containing the sink node (i.e., sink stage) uses the $k_{b}$$k_{b}$ schedule (we introduces GraphPipe’s micro-batch schedules in §6). These supposed conditions comprise a schedule configuration denoted by $c:=(i,b,k)$ in Algorithm 2. They are either available as the results of some other DP subproblems solved previously, or furnished by enumeration. The solution of this DP subproblem computes the smallest possible number of in-flight micro-batches for the source stage (i.e., $i_{f}$ in Figure 4) that meets the target TPS range $[0,t_{max}]$.
@@
6. Static Micro-Batch Scheduler
The static micro-batch scheduler of GraphPipe optimizes micro-batch schedules to minimize training time and memory footprint. Specifically, we design our scheduler to address the unique challenges presented by graph-like data dependencies in GPP pipeline stages. These dependencies complicate both scheduling and memory footprint optimization. Our scheduler aims to optimize schedules while ensuring their validity by meeting the data dependency requirements between all forward and backward passes.
We further generalize our scheduler so that it can support different micro-batch sizes and schedules over pipeline stages. This can be effective when running heterogeneous models (e.g., multi-modal models with different ideal micro-batch size over pipeline stages across different modalities). We also find it can save extra memory footprint with fine-grained per-stage micro-batch scheduling (i.e., Figure 5).
The scheduler (Algorithm 2) takes as input (1) a configuration of model partition $\mathcal{G}$, (2) current and next stage schedule configurations $c_{f},c_{b}$, and (3) the number of devices $d$ from the pipeline stage partitioner, and produces an optimized micro-batch schedule $\Pi_{opt}$ for a given stage configuration. As in Figure 3, the input is fed by the stage partitioner, and the output is returned back to the stage partitioner to form a stage graph with an optimized micro-batch schedule.
@@ 6. Static Micro-Batch Scheduler
[…]
@@
Second, $\textsc{ScheduleTask}(\cdot)$ produces an optimized schedule of forward and backward passes for a given stage with optimized schedule configuration ($c_{opt}$). It adopts greedy scheduling that schedules backward pass as early as possible. It reduces both memory consumption and training iteration time since it quickly resolves the corresponding in-flight forward pass. However, graphical pipeline stages and different micro-batch sizes over stages make the data dependencies of forward and backward passes convoluted. Still, we find a greedy scheduling handling these factors is optimal in terms of iteration time and memory footprint.
@@ 7. Evaluation
We develop GraphPipe on top of FlexFlow (jia2019beyond, ), a distributed multi-GPU runtime for DNN training. Major modifications are to replace FlexFlow’s partitioner and scheduler with ours described in §5 and §6 respectively. We evaluate GraphPipe on the Summit supercomputer (summit, ). For each compute node of Summit, we use 2 IBM POWER9 CPUs and 4 NVIDIA V100 GPUs with 512GB of main memory. GPUs within a node are interconnected via NVLink while nodes are connected via Mellanox EDR 100Gb InfiniBand. We use the default schedule configurations of GraphPipe mentioned in Section 6. Note that we omit error bars for our plots, as we observe marginal standard deviations (less than 3%) for all results.
DNNs. We explore three multi-branch DNNs: Multi-Modal Transformer-based model (MMT) (vaswani2017attention, ; radford2021learning, ), DLRM (naumov2019deep, ), and CANDLE-Uno (2018candleuno, ). Multi-Modal Transformer (MMT) is a backbone of most state-of-the-art multi-modal models (wang2023large, ; radford2021learning, ; openai2023gpt4, ; ramesh2021zero, ; jia2021scaling, ). DLRM is a popular deep learning recommendation model for personalization and ads recommendation. CANDLE-Uno is a specialized model in the medical domain (i.e., precision medicine). We describe the detailed model configurations in the supplemental material. Despite different applications, all these models feature parallel branches, each processing a different type of data.
@@ 7.1. End-to-End Evaluation
We compare the training throughput of GraphPipe with existing pipeline-parallel systems: PipeDream (narayanan2021memory, ) and Piper (tarnawski2021piper, ). We choose these two baselines since their combined search space encompasses all possible model partitions covered by other approaches including DAPPLE (fan2021dapple, ), Alpa (zheng2022alpa, ), etc (narayanan2019pipedream, ; unger2022unity, ). We implemented their stage partitioning strategies in GraphPipe to conduct fair comparisons; note that it is also necessary for Piper given that it is a partitioning algorithm and doesn’t have a runtime.
@@
We compare the training throughput of GraphPipe with existing pipeline-parallel systems such as PipeDream (narayanan2021memory, ) and Piper (tarnawski2021piper, ). We choose these two baselines since their combined search space encompasses all possible model partitions covered by other approaches (fan2021dapple, ; zheng2022alpa, ; narayanan2019pipedream, ).
@@ 7.4. Case Study
It is instructive to take a close look at the strategies produced by GraphPipe and SPP. We run both GraphPipe and SPP optimizers for a synthetic model, execute the strategies, and observe a 20% throughput improvement by GraphPipe over SPP. Our analysis finds that the aggregate gain comes from two sources, and the contributions are nearly equal.

### v2

@@ 1. Introduction
First, opportunities to exploit the inherent parallel structures of a DNN are left unseized. DNN applications such as healthcare (krumholz2016data, ; tan2020multimodal, ; suresh2020deep, ), chatbot (openai2023gpt4, ), and recommendation (naumov2019deep, ) jointly process heterogeneous data types (e.g., text, images, and tabular data). Specifically, the rise of generalist AI models, such as GPT-4o (gpt-4o, ), Chameleon (team2024chameleon, ), and Gato (reed2022generalist, ), further underscores the need for efficient parallel handling of diverse data modalities. DNNs employed therein are designed to feature multiple branches, which are computationally independent and thus can be executed concurrently. But existing DNN systems with SPP first linearize the computation graph of a DNN to construct the stages of a sequential pipeline and process these stages sequentially, falling short in harnessing the opportunity to blend such branch-level parallelism with pipeline parallelism.
@@ 5. Pipeline Stage Partitioner
22:// Dynamic Programming (DP) Partitioner
23:function DP($\mathcal{G},c_{f},c_{b},d,t_{max}$)
24:    if this DP state has been visited then
25:         return corresponding $\mathcal{G}_{S}^{best}$ to this DP state
26:    // Consider a given DP state as a SINGLE stage
27:    $\mathcal{G}_{S}^{best}=\varnothing$
28:    if EstimateTPS($\mathcal{G},c_{f},c_{b},d$)$\leq t_{max}$ then
29:         // Optimize schedule via Algorithm 2
30:         $\Pi_{opt}$ = ScheduleStage($\mathcal{G},c_{f},c_{b},d$)
31:         $\mathcal{G}_{S}^{best}=$ StageGraph($\mathcal{G},\Pi_{opt},d$)
32:    // Decompose a given DP state into two stages
33:    if $\mathcal{G}$ can be decomposed in series then
34:         for $(\mathcal{G}_{1},\mathcal{G}_{2})\in$ SeriesDecompose($\mathcal{G}$do
35:             for $d_{2}\leftarrow 1$ to $d-1$ do
36:                 $d_{1}=d-d_{2}$
37:                 for $c_{m}\in C$ do
38:                     $\mathcal{G}_{S_{2}}^{new}=$ DP($\mathcal{G}_{2},c_{m},c_{b},d_{2},t_{max}$)
39:                     Update $i_{m}$ based on $\mathcal{G}_{S_{2}}^{new}$
40:                     $\mathcal{G}_{S_{1}}^{new}=$DP($\mathcal{G}_{1},c_{f},c_{m},d_{1},t_{max}$)
41:    else if $\mathcal{G}$ can be decomposed in parallel then
42:         for $(\mathcal{G}_{1},\mathcal{G}_{2})\in$ ParallelDecompose($\mathcal{G}$do
43:             for $d_{1}\leftarrow 1$ to $d-1$ do
44:                 $d_{2}=d-d_{1}$
45:                 $\mathcal{G}_{S_{1}}^{new}=$ DP($\mathcal{G}_{1},c_{f},c_{b},d_{1},t_{max}$)
46:                 $\mathcal{G}_{S_{2}}^{new}=$ DP($\mathcal{G}_{2},c_{f},c_{b},d_{2},t_{max}$)
47:    $\mathcal{G}_{S}^{best}=$ PickBetter($\mathcal{G}_{S}^{best},\mathcal{G}_{S_{1}}^{new}\cup\mathcal{G}_{S_{2}}^{new}$)
48:    return $\mathcal{G}_{S}^{best}$
@@
Dynamic programming (DP). GraphPipe adopts a dynamic programming algorithm where the value of each DP state indicates the existence of a strategy achieving a throughput within a target range (Lines 1320 of Algorithm 1). At each DP level, GraphPipe applies series-parallel decompositions to split an input graph (say $\mathcal{G}$) into two new subgraphs (say $\mathcal{G}_{1},\mathcal{G}_{2}$), each of which serves as the input computation graph of a new DP subproblem at one DP level below. GraphPipe recursively solves the DP subproblems to construct a solution of the original problem where the input computation graph is $\mathcal{G}_{C}$ (Lines 2348 of Algorithm 1).
@@
  • Base case: We consider the entire subgraph $\mathcal{G}$ as a single stage and apply data parallelism with a data-parallel degree of $d$ (Line 28 in Algorithm 1). We estimate TPS by profiling the execution time of each operator while extrapolating communication latency by affine functions. We check if the target TPS range is achievable with the memory constraint, and compute the number of in-flight micro-batches according to Algorithm 2 (see §6).
  • Series decomposition: We perform a series decomposition to create two subgraphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, where the sink node of $\mathcal{G}_{1}$ coincides with the source node of $\mathcal{G}_{2}$ (Line 33 in Algorithm 1). We first solve the subproblem associated with $\mathcal{G}_{2}$. To do so, we enumerate all feasible schedules for the source node of $\mathcal{G}_{2}$. We then solve the subproblem associated with $\mathcal{G}_{1}$.
  • Parallel decomposition: We perform a parallel decomposition to create $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, where $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ share the same source and sink nodes (Line 41 in Algorithm 1). As there is no data dependency between these subgraphs, the pipelines can be executed in parallel. The subproblems associated with $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ may produce different optimal numbers of in-flight micro-batches for the shared source node. To ensure continuous pipelining, we take the larger number of in-flight micro-batches as the solution.
Overall process. Figure 4 visualizes the overall process. At the top, a DP subproblem is provided with its initial conditions: computation graph $\mathcal{G}$, the number of available devices $d$, and the target TPS range $[0,t_{max}]$. Suppose the number of in-flight micro-batches for the sink node is $i_{b}$, the micro-batch sizes for the source and sink nodes are $b_{f}$ and $b_{b}$, the stage containing the source node (i.e., source stage) uses the $k_{f}$$k_{f}$ schedule, and the stage containing the sink node (i.e., sink stage) uses the $k_{b}$$k_{b}$ schedule (we introduces GraphPipe’s micro-batch schedules in §6). These supposed conditions comprise a schedule configuration denoted by $c:=(i,b,k)$ [^1] Find the definition of $i,b,k$ in Section A.1 in Algorithm 2. They are either available as the results of some other DP subproblems solved previously, or furnished by enumeration. The solution of this DP subproblem computes the smallest possible number of in-flight micro-batches for the source stage (i.e., $i_{f}$ in Figure 4) that meets the target TPS range $[0,t_{max}]$.
@@
Figure 5. A comparison between universal and per-stage micro-batch size / schedule. F$\{i,j\}$, B$\{i,j\}$ indicate forward and backward passes for a micro-batch including samples $i$ and $j$. It showcases how per-stage micro-batch size and scheduling can save memory footprint and training iteration time.
6. Static Micro-Batch Scheduler
[…]
We further generalize our scheduler so that it can support different micro-batch sizes and schedules over pipeline stages. This can be effective when running heterogeneous models (e.g., multi-modal models with different ideal micro-batch size over pipeline stages across different modalities).
Figure 5 illustrates how we can reduce training iteration time and memory footprint with per-stage micro-batch size and scheduling. Here, each pipeline stage has different smallest micro-batch size (i.e., 1, 2, 4 for $S_{1},S_{2},S_{3}$) achieving maximum compute efficiency. Using a fixed micro-batch size of 4 across all stages maximizes compute efficiency. But, the drawback is long GPU idling during warm-up and cool-down, and large memory footprint, i.e., 12 in-flight micro-batches for stage 1. However, by tailoring the micro-batch size and scheduling to each stage, we reduce this to 10 in-flight micro-batches for stage 1 while maintaining maximum compute efficiency. This also shortens the training iteration because stages can be scheduled earlier with smaller micro-batch sizes. This benefit usually grows with more pipeline stages.
@@ 6. Static Micro-Batch Scheduler
To support 1) graph-like stage dependency and 2) per-stage micro-batch size and schedule, this is how we design our scheduler (Algorithm 2). It takes as input (1) a configuration of model partition $\mathcal{G}$, (2) current and next stage schedule configurations $c_{f},c_{b}$, and (3) the number of devices $d$ from the pipeline stage partitioner, and produces an optimized micro-batch schedule $\Pi_{opt}$ for a given stage configuration. As in Figure 3, the input is fed by the stage partitioner, and the output is returned back to the stage partitioner to form a stage graph with an optimized micro-batch schedule.
@@
$\textsc{ComputeInFlight}(\cdot)$ and $\textsc{ScheduleTask}(\cdot)$ are two key functions in Algorithm 2. First, $\textsc{ComputeInFlight}(\cdot)$ is a key subroutine to optimize schedule by effectively minimizing the number of in-flight micro-batches for a given stage without increasing per-iteration training time. To take graphical stage dependency (i.e., multiple stages following one) into account, it factors in all following stages to decide the minimum number of in-flight micro-batches. It also accounts for scheduling constraints posed by micro-batch size gap between subsequent stages. For instance, in Figure 5, $S_{2}$ needs two micro-batches to be processed from $S_{1}$ to process a single micro-batch. Section A.1 explains the detail of $\textsc{ComputeInFlight}(\cdot)$ computes the minimal number of in-flight micro-batches.
@@
Second, $\textsc{ScheduleTask}(\cdot)$ produces an optimized schedule of forward and backward passes with optimized schedule configuration ($c_{opt}$). It adopts greedy scheduling that schedules backward pass as early as possible. It reduces both memory consumption and training iteration time since it quickly resolves the corresponding in-flight forward pass.
@@ 7. Evaluation
We develop GraphPipe on top of FlexFlow (jia2019beyond, ), a distributed multi-GPU runtime for DNN training. We adjusted FlexFlow’s runtime for parallel execution of graphical pipeline stages while introducing our own partitioner in §5 and scheduler in §6 to FlexFlow. We evaluate GraphPipe on the Summit supercomputer (summit, ). For each compute node of Summit, we use 2 IBM POWER9 CPUs and 4 NVIDIA V100 GPUs with 512GB of main memory. GPUs within a node are interconnected via NVLink while nodes are connected via Mellanox EDR 100Gb InfiniBand. We use the default schedule configurations of GraphPipe mentioned in Section 6. Note that we omit error bars for our plots, as we observe marginal standard deviations (less than 3%) for all results.
DNNs. We explore three multi-branch DNNs: Multi-Modal Transformer-based model (MMT) (vaswani2017attention, ; radford2021learning, ), DLRM (naumov2019deep, ), and CANDLE-Uno (2018candleuno, ). Multi-Modal Transformer (MMT) is a backbone of most state-of-the-art multi-modal models (wang2023large, ; radford2021learning, ; openai2023gpt4, ; ramesh2021zero, ; jia2021scaling, ). DLRM is a popular deep learning recommendation model for personalization and ads recommendation. CANDLE-Uno is a specialized model in the medical domain (i.e., precision medicine). We describe the detailed model configurations in Section A.2. Despite different applications, all these models feature parallel branches, each processing a different type of data.
@@ 7.1. End-to-End Evaluation
We compare the training throughput of GraphPipe with existing pipeline-parallel systems such as PipeDream (narayanan2021memory, ) and Piper (tarnawski2021piper, ). We choose these two baselines since their combined search space encompasses all possible model partitions covered by other SPP approaches (fan2021dapple, ; zheng2022alpa, ; narayanan2019pipedream, ). To be specific, PipeDream (with the operator granularity) basically covers the pipeline partitioning and scheduling strategies of all baseline SPP approaches (fan2021dapple, ; zheng2022alpa, ; narayanan2019pipedream, ) but Piper. They all (1) linearize DNNs by transforming their computation graphs into sequences of operators, (2) exhaust pipeline stage partition choices, and (3) employ 1F1B scheduling.
@@ 7.3. Different Numbers of Branches and
Figure 9. Ablation study: end-to-end throughput comparison of different strategies on different models.
@@
7.4. Ablation Study
Figure 9 shows the breakdown of performance benefits of GraphPipe from 1) parallel execution of stages and 2) increased micro-batch size from reduced memory footprint. In Figure 9, ”Parallel” is the strategy that allows parallel execution of stages with same micro-batch size with SPP while ”GraphPipe” is the strategy that allows both parallel execution of stages and larger micro-batch size than SPP. Note that It is not possible to evaluate the strategy only with larger micro-batch size since the reduced pipeline depth from parallel stage execution enables larger micro-batch size. We evaluate throughput of each strategy with 32 GPUs.
We observe that, compared to SPP, ”Parallel” strategy achieves 1.12 - 1.40$\times$ speedup while ”GraphPipe” strategy achieves 1.25 - 1.61$\times$ speedup. This result indicates that both performance benefits of GraphPipe are crucial. We also find the consistent pattern of the strategy optimized by GraphPipe across different models. In the following section, we explain how these two sources of performance gains contribute to the overall improvement achieved by GraphPipe.
@@ 7.5. Case Study
To clearly illustrate the advantages of GraphPipe, we analyze the strategies it produces in comparison to SPP, using a simplified synthetic model for clarity. We run both GraphPipe and SPP optimizers, execute the strategies, and observe a 20% throughput improvement by GraphPipe over SPP. Our analysis finds that the aggregate gain comes from two sources, and the contributions are nearly equal.
@@
GraphPipe’s performance improvement becomes more significant as 1) memory pressure and 2) the parallelism degree within a DNN increase. The selection of hardware influences memory pressure, subsequently affecting GraphPipe’s performance improvement, while it is common practice for the system to operate close to memory limits in DNN training. In contrast, the parallelism degree within a DNN is hardware independent, and therefore GraphPipe’s performance improvement over existing systems will maintain across different hardware platforms.
@@ 9. Conclusion
Acknowledgement
We would like to thank members of Catalyst group at CMU for their helpful comments on our work and manuscript. We would also like to thank the anonymous ASPLOS reviewers for constructive feedbacks. This work was partially supported by the National Science Foundation under grant numbers CNS-2147909, CNS-2211882, and CNS-2239351, along with gift awards from Amazon, Cisco, Google, Meta, Oracle, Qualcomm, and Samsung. Additional support was provided by the Real Time Machine Learning (RTML) DARPA project.
Appendix A Appendix
[…]
The $k_{x}$$k_{x}$ schedule of stage $S_{x}$ is determined by
$\mathrm{argmin}_{k_{x}}\max_{(S_{x},S_{y})\in\mathcal{V}_{S}}\mathrm{ComputeInFlight}(k_{x},b_{x},k_{y},b_{y},i_{y}),$
where $i_{y},b_{y}$ are the number of in-flight samples and micro-batch size for stage $S_{y}$. $\mathrm{ComputeInFlight}(k_{x},b_{x},k_{y},b_{y},i_{y})$ is computed according to Table 2:
Condition Result
$\max\{b_{x},b_{y}\}<k_{x}b_{x}<k_{y}b_{y}$ $i_{y}+2\max\{b_{x},b_{y}\}$
$\max\{b_{x},b_{y}\}=k_{x}b_{x}<k_{y}b_{y}$ $i_{y}+\max\{b_{x},b_{y}\}$
$b_{x}\leq b_{y}<k_{y}b_{y}<k_{x}b_{x}$ $i_{y}+k_{x}b_{x}-k_{y}b_{y}+2b_{y}$
$b_{x}\leq b_{y}=k_{y}b_{y}<k_{x}b_{x}$ $i_{y}+k_{x}b_{x}$
$b_{y}\leq b_{x}<k_{y}b_{y}<k_{x}b_{x}$ $i_{y}+k_{x}b_{x}-k_{y}b_{y}+2b_{x}$
$b_{y}\leq b_{x}=k_{y}b_{y}<k_{x}b_{x}$ $i_{y}+k_{x}b_{x}$
$\max\{b_{x},b_{y}\}=k_{y}b_{y}=k_{x}b_{x}$ $i_{y}+k_{y}b_{y}$
$\max\{b_{x},b_{y}\}<k_{y}b_{y}=k_{x}b_{x}$ $i_{y}+2\max\{b_{x},b_{y}\}$
$b_{x}\leq k_{x}b_{x}<b_{y}\leq k_{y}b_{y}$ $i_{y}+b_{y}$
$b_{y}\leq k_{y}b_{y}<b_{x}\leq k_{x}b_{x}$ $i_{y}+k_{x}b_{x}-k_{y}b_{y}+b_{x}$
Table 2. Computation of the number of in-flight samples.
A.2. DNN Model Configurations
The Multi-Modal Transformer-based model (MMT) for which we evaluate GraphPipe consists of four parallel branches concatenated at the end and each branch consists of eight Transformer layers (32 layers in total). Here, the input sequence length is 256. Each transformer layer has a hidden size of 1024, an embedding size of 1024, and 16 attention heads. The hidden size for a feed-forward layer following the attention layer has a hidden size of 4096.
The DLRM model for which we evaluate GraphPipe consists of seven branches for dense features and seven branches for sparse features (embedding layers); these branches are concatenated at the end. Each branch for dense features includes four feed-forward layers. The hidden size of dense features and the following feed-forward layers is 4096. For sparse features, its hidden size is 64 and the embedding bag size is 100; embeddings in a single bag is concatenated. The number of entries in an embedding table is 1 million. Feed-forward layers post-processing the interaction also have the hidden size of 4096.
The CANDLE-Uno model for which we evaluate GraphPipe consists of seven branches, each of which includes four feed-forward layers. All feed-forward layers have a hidden size of 4096.
For our end-to-end evaluations, we use the following ranges of mini-batch sizes for each device count such that the system operates close to the memory limit:
# Devices MMT DLRM CANDLE-Uno
4 64 256 4096
8 128 512 8192
16 256 1024 16384
32 512 2048 32768
We sweep over all possible micro-batch sizes given mini-batch sizes for each model to maximize training throughput.
A.3. Sequential DNN Evaluation
We evaluate GraphPipe with baselines and confirm that GraphPipe match performance of baselines when the workload is sequential DNN. We measure throughput on the sequential Transformer model with the same configuration with the MMT model above.
# Devices Piper PipeDream Ours
4 113.94 113.2 113.6
8 187.92 190.1 189.5
16 286.83 277.00 287.52
32 609.76 592.4 608.82